OpenStudy (anonymous):

the equation ax-4y+2=0 represents a linear relation. Determine the value of A if the point (6,5) belongs

OpenStudy (anonymous):

(6, 5) can be said to give us an x-value of 6 and a y-value of 5 ax-4y+2=0 a(6) - 4(5) + 2 = 0 6a - 20 + 2 = 0 6a - 18 = 0 6a = 18 a = 3